OpenStudy (anonymous):

Calc 3 - Matching quadric surfaces to equations??

OpenStudy (anonymous):

I can't seem to get any of them right? How would you know what matches with what.

OpenStudy (anonymous):

I took each equation and divided and moved them around until the right side had a one

OpenStudy (turingtest):

well g is the sphere because it's just the 3D version of ax^2+bx^2=r^2

OpenStudy (anonymous):

Oh I thought that a b and c had to be 1 to be a perfect sphere

OpenStudy (turingtest):

http://tutorial.math.lamar.edu/Classes/CalcII/QuadricSurfaces.aspx i forgot the rest but this page is great for this stuff

OpenStudy (anonymous):

do you know what the second one is suppose to be?

OpenStudy (anonymous):

I can't even identify the shape, is it a plane?

OpenStudy (turingtest):

right, sorry, it's not a sphere but a ellipsoid I have no idea what that second shape is either :( check out that page I linked you to though.

OpenStudy (anonymous):

I already had a look at that previous it's confusing b/c it gives a general equation of the shape

OpenStudy (anonymous):

but thanks for your help

OpenStudy (turingtest):

the cone is D because you can write it as X^2+Y^2=Z^2 forget the coeffifents and focus on the signs and degrees of the variables

OpenStudy (turingtest):

shape A is eqn. H because it can be written as x^2+y^2=z^2+N

OpenStudy (anonymous):

I think.. your first figure is not sphere it is elsiod and it's matches with G your 2nd figure is plane and it matches with B your 3rd figure is hyperboliod and it matches with A your 4th figure is cone and it matches with G

OpenStudy (turingtest):

you're right about the first three, but the cone has to be D (it can't be G again, could it?) look at the form and compare it to the one on the link

OpenStudy (anonymous):

you r ri8 I wanted to write D but by mistake I wrote G sorry..

OpenStudy (anonymous):

Whoa you are right, how did you know the third figure was a hyperboliod?

OpenStudy (anonymous):

at plane x=0 you can see it's hyperbola